OpenStudy (anonymous):

Which of the following is an equation of the line that passes through the point (1, 1) and has a slope of 2? A. y = 2x – 1 B. y = 2x + 1 C. y = –2x – 1 D. y = –2x + 1

undeadknight26 (undeadknight26):

let see.. y = mx + b right? m = slope and b=y

undeadknight26 (undeadknight26):

I believe it is B.

OpenStudy (texaschic101):

y = mx + b slope(m) = 2 (1,1) x = 1 and y = 1 now we sub.. 1 = 2(1) + b 1 = 2 + b 1 - 2 = b -1 = b so you equation is : y = 2x - 1